OpenStudy (notcheatinglol):

Find the exact value of : cos(-7pi/4)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got a unit circle cheat sheet?

OpenStudy (notcheatinglol):

nope, homeschooled

OpenStudy (notcheatinglol):

i like to learn as i take the test, i learn alot better that way

OpenStudy (ooxbrittanyxoo):

\[\sqrt{2}/2\]

OpenStudy (notcheatinglol):

^ can you walk me though it?

OpenStudy (misty1212):

find \(-\frac{7\pi}{4}\) on the unit circle on the last page, then look at the corresponding point the first coordinate is cosine, the second coordinate is sine

OpenStudy (misty1212):



OpenStudy (misty1212):

\[-\frac{7\pi}{4}\] means count out clockwise in unit of \(\frac{\pi}{4}\)

OpenStudy (notcheatinglol):

ohhhhh! thanks misty<3 Nice and pretty haha i may have some more question sbut i got this now!

OpenStudy (misty1212):



OpenStudy (notcheatinglol):

i think i is\[\sqrt{2}\]

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

don't forget the denominator

OpenStudy (notcheatinglol):

its 2 right

OpenStudy (misty1212):

yes 2 is the denominator, answer is \[\frac{\sqrt2}{2}\]